# Canonical decompositions of 3-connected graphs

### v2

@@ 1.3. Nested or crossed: analysing corner diagrams
Suppose that $(C,D)$ is a tri-separation of $G$ that crosses $(A,B)$. If $(C,D)$ is strong, we are done, so we may assume that some vertex $u$ of $G$ of degree three lies in the separator of $(C,D)$. Since $(A,B)$ is a strong tri-separation, the vertex $u$ cannot lie in the centre, so $u$ lies in a link, say it lies in the link for $A$. As a $K_{4}$ has no strong nontrivial tri-separation, $G$ is not a $K_{4}$. By Section 1.2, $(C,D)$ is nontrivial. Hence we may apply the Nested or crossed: analysing corner diagrams (1.3) to find that the existence of $u$ implies that all links have size one while the centre consists of a single vertex, and that there are no jumping edges. Since $(C,D)$ is a tri-separation and $u$ has degree three, $u$ has a neighbour $v$ in $C\cap D$. As there are no jumping edges, the neighbour $v$ of $u$ can only lie in the centre. Since $(A,B)$ is a tri-separation, $v$ has a neighbour $w$ in $A$ besides $u$. By symmetry $w\in C$.
Claim 1.3.0.2.
The vertex $v$ has three neighbours in $C$.
Proof of Claim. If the corner $\{B,C\}$ is nonempty, then as $G$ is 3-connected the corner contains a neighbour of $v$, and together with $u$ and $w$ we have found three neighbours of $v$ in $C$. Thus assume that the corner $\{B,C\}$ is empty. Since not both adjacent links can consist of an edge, at least one adjacent link contains a vertex $y$. Since the corner for $\{B,C\}$ is empty but $y$ has two neighbours either in $B$ or in $C$ (depending on which tri-separator $S(A,B)$ or $S(C,D)$ contains $y$), it follows that $y$ is adjacent to $v$. If $y$ is in the link for $B$, then $u$, $y$ and $w$ are three distinct neighbours of $v$ in $C$, and we were done. So assume that $y$ is in the separator of the strong tri-separation $(A,B)$. Thus $y$ has degree at least four. And since the corner $\{B,C\}$ is empty, the vertex $y$ must have one of its neighbours outside the mixed 3-separator $S(C,D)$ in the corner $\{A,C\}$, and this nonempty corner contains a neighbour of $v$ by 3-connectivity, which is different from $u$ and $y$. $\diamondsuit$
@@
Let $C^{\prime}:=C-u$ and $D^{\prime}:=D$. Then the separator of $(C^{\prime},D^{\prime})$ arises from the separator of $(C,D)$ by replacing the vertex $u$ with the edge $uc$. The only vertex in the separator of $(C,D)$ that might loose a neighbour when moving to $(C^{\prime},D^{\prime})$ is the vertex $v$, which looses its neighbour $u$ in $C^{\prime}$. However, Claim 1.3.0.2 ensures that $v$ has two neighbours in $C^{\prime}$. Hence $(C^{\prime},D^{\prime})$ is a tri-separation, and it has less vertices of degree three in its separator than $(C,D)$. Moreover, $(C^{\prime},D^{\prime})$ crosses $(A,B)$ with the same links and centre as for $(C,D)$, with just one exception: the link for $A$, which consisted of $u$ for $(C,D)$, consists of the edge $uc$ for $(C^{\prime},D^{\prime})$. By iterating at most two times, we obtain a strong tri-separation that crosses $(A,B)$. ∎
@@
(1$\to$(2). We only have to show that $(A,B)$ is strong. This follows from Section 1.3.
(2$\to$(1). Suppose for a contradiction that $(A,B)$ is crossed by a tri-separation $(C,D)$ of $G$. By Section 1.3, we may assume that $(C,D)$ is strong. Since $(A,B)$ is strong, $(C,D)$ is nontrivial by Section 1.2. This contradicts (1). ∎
[…]
Proof.
By Section 1.2, $(A,B)$ is nontrivial. Since $S(A,B)$ consists of edges, $(A,B)$ is strong. By Section 1.3, it suffices to show that $(A,B)$ is nested with every strong nontrivial tri-separation of $G$. And indeed, since $S(A,B)$ contains no vertices and since $K_{4}$ has no non-atomic 3-cut, $(A,B)$ is nested with every strong nontrivial tri-separation of $G$ by the Nested or crossed: analysing corner diagrams (1.3). ∎
1.4. Proof of Applications
Before we prove the (Angry Tri-Separation Theorem)., let us see how it implies Applications. A graph $G$ is essentially 4-connected if it is 3-connected, every nontrivial strong tri-separation has three edges in its separator such that the subgraph induced by one side is equal to a triangle, and $G\neq K_{4}$. A graph $G$ is vertex-transitive if the automorphism group of $G$ acts transitively on its vertex set $V(G)$.
@@ 1.4. Proof of Applications
Let $G$ be a vertex-transitive finite connected graph. We have to show that $G$ either is essentially 4-connected, a cycle, or a complete graph on at most four vertices. By Section 3.3.3, $G$ is a cycle, $K_{2}$, $K_{1}$ or 3-connected. Since we are done otherwise, let us assume that $G$ is 3-connected. By the (Angry Tri-Separation Theorem). (1.1.1), $G$ is internally 4-connected, a $K_{3,m}$ with $m\geqslant 3$, a wheel, or $G$ has a totally-nested nontrivial tri-separation. If $G$ is internally 4-connected, then $G\notin\{K_{4},K_{3,3}\}$ by definition, and all strong tri-separations of $G$ are trivial by Section 1.2; in particular, $G$ is essentially 4-connected. If $G$ is a $K_{3,m}$ for some $m\geqslant 3$, then $m=3$ since $G$ is vertex-transitive, and $G=K_{3,3}$ is essentially 4-connected since all its strong tri-separations are trivial. If $G$ is a wheel, then $G$ can only be a $K_{4}$ by vertex-transitivity, and $K_{4}$ is a possible outcome.
@@
Proof of Claim. Suppose for a contradiction that there is a vertex $v\in U\cap W$. Since $(U,W)$ is a mixed-separation, there is a vertex $u\in U\setminus W$. Let $\varphi\in\text{Aut}(G)$ send $v$ to $u$. Then $(\varphi(U),\varphi(W))\leqslant(U,W)$ or $(\varphi(W),\varphi(U))\leqslant(U,W)$ since $(U,W)$ is totally-nested, and in either case the inequality is strict since $u$ does not lie in $S(U,W)$ but does so after the application of $\varphi$. This contradicts the choice of $(U,W)$. $\diamondsuit$
Claim 1.4.0.2.
@@
Proof of Claim. Since $G$ is 3-connected, since $(U,W)$ is nontrivial and since $S(U,W)$ consists of three edges by Claim 1.4.0.1, it suffices to show that every vertex in $U$ is incident with an edge in $S(U,W)$. The proof is analogue to the proof of Claim 1.4.0.1. $\diamondsuit$
By Claim 1.4.0.1 and Claim 1.4.0.2, every tri-separation in $O$ has three edges in its separator and a side that induces a triangle. As $(A,B)$ was chosen arbitrarily, every totally-nested nontrivial tri-separation has three edges in its separator and a side that induces a triangle. Hence to show that every nontrivial strong tri-separation also has three edges in its separator and a side that induces a triangle, it suffices to show that
Claim 1.4.0.3.
@@
Proof of Claim. Since $S(U,W)$ consists of three edges (which share no ends by Section 1.2) and $G[U]=K_{3}$, all vertices of $G$ have degree three. Hence all vertices of $G$ have degree three. Let $(C,D)$ be an arbitrary nontrivial strong tri-separation of $G$. As $(C,D)$ is strong, the separator of $(C,D)$ consists of three edges. Then $(C,D)$ is totally-nested by Section 1.3. $\diamondsuit$
Combining Claim 1.4.0.1, Claim 1.4.0.2 and Claim 1.4.0.3 yields that $G$ is essentially 4-connected. ∎
@@
1.5. Understanding nestedness through connectivity
@@
Figure 1.5. The situation excluded by Section 1.5
For a depiction of the situation excluded by our next lemma, see Figure 1.5.
@@
Let $(C,D)$ be a strong tri-separation of $G$ that crosses $(A,B)$ so that $x_{3}$ is in the centre. Since $K_{4}$ has no strong tri-separation, $G$ cannot be a $K_{4}$. By the Nested or crossed: analysing corner diagrams (1.3) and since $(A,B)$ is half-connected, $x_{3}$ is a vertex and the only element of the centre, all links have size one, and there are no jumping edges. Without loss of generality, $x_{2}$ lies in the link for $C$, and $y\in A\setminus B$. If $y$ lies in the corner for $\{A,C\}$, then we are done. Otherwise, $y$ must lie in the link for $A$. The corner for $\{A,C\}$ must be empty, since otherwise $\{y,x_{3}\}$ would be a 2-separator of $G$, contradicting 3-connectivity. As $y\in S(C,D)$ must have two neighbours in $G[C]$, and since there are no jumping edges, it follows that $yx_{3}$ must be an edge in $G$. But $y\in S(C,D)$ also means that $y$ cannot have degree three as $(C,D)$ is strong, and since $y$ is $x_{3}$-free this means that the edge $yx_{3}$ must not be present in $G$, a contradiction. ∎
Figure 1.6. The situation in the second and third case of the proof of Section 1.5
@@
Let us denote the separator of $(A,B)$ by $\{x_{1},x_{2},x_{3}\}$, and let us assume that $x_{3}$ is a vertex and that the separator is externally tri-connected around $x_{3}$. Let us assume for a contradiction that $(A,B)$ is crossed by a strong tri-separation $(C,D)$ so that $x_{3}$ lies in the centre. Since $K_{4}$ has no strong tri-separation, $G$ is not a $K_{4}$. By the Nested or crossed: analysing corner diagrams (1.3), all links have size one, $x_{3}$ is the only element of the centre, and there are no jumping edges. We distinguish three cases.
@@
($=$) In the third case, $x_{1}$ and $x_{2}$ are edges. By external tri-connectivity, these edges have $x_{3}$-free endvertices $y_{1}$ and $y_{2}$ and there are two internally disjoint paths $P,Q$ from $y_{1}$ to $y_{2}$ avoiding $x_{1},x_{2},x_{3}$. By symmetry assume that the vertex $y_{1}$ lies in the side $A$. Then the two paths $P,Q$ are contained in $G[A]$ and $y_{2}$ is in $A$ as well. Since the link for $A$ has size one, not both of the two paths $P,Q$ can meet it in internal vertices. So $y_{1}$ or $y_{2}$ must lie in the link for $A$. This contradicts Section 1.5. ∎
@@
Let $(A,B)$ be a tri-separation of $G$ that is externally tri-connected, half-connected, strong and nontrivial. Assume for a contradiction that $(A,B)$ is crossed by a tri-separation $(C,D)$ of $G$. By Section 1.3, we may assume that $(C,D)$ is strong and nontrivial. As $K_{4}$ has no strong nontrivial tri-separation, $G$ is not a $K_{4}$. Hence by the Nested or crossed: analysing corner diagrams (1.3), $(A,B)$ and $(C,D)$ cross so that the centre contains a vertex. This contradicts Section 1.5. ∎
@@
Proof of Section 1.5.
Note first that $E(A\setminus B,B\setminus A)$ is empty since $G$ is 3-connected. Hence it suffices to find three internally disjoint paths between any pair of vertices in the separator $A\cap B$ avoiding the third vertex, by criterion (:) and Section 1.5. By assumption, $G\setminus(A\cap B)$ has at least three components, and each component has neighbourhood equal to $A\cap B$ since $G$ is 3-connected. Thus we find three internally disjoint paths for each pair of vertices in $A\cap B$ through these components. ∎
@@
(1). Since $G[B]$ has minimum degree two, we deduce that $B^{\prime}=B$, and so $(A^{\prime},B^{\prime})\leqslant(A,B)$.
@@
(4). If $(A^{\prime},B^{\prime})$ is nontrivial, then $G[A]$ contains a cycle, so two vertices in $X$ are adjacent or $|K|\geqslant 2$. Conversely, suppose now that two vertices in $X$ are adjacent or that $|K|\geqslant 2$. Since $B^{\prime}=B$ and $|B|\geqslant 2$, it follows that $|B^{\prime}|\geqslant 2$. Thus it suffices to show that $A^{\prime}$ contains at least two vertices, by Section 1.2. If two vertices in $X$ are adjacent, then these two vertices are not reduced to edges in $S(A^{\prime},B^{\prime})$, so they lie in $A^{\prime}$ and we are done. So assume that $|V(K)|\geqslant 2$. Since $V(K)\subseteq A^{\prime}$, we are done as well. ∎
@@
As $(C,D)$ is not half-connected, by 3-connectivity of $G$ its separator $X$ consists of three vertices; and $G\setminus X$ has at least four components. By Section 1.5, no two vertices in $X$ are adjacent, and every component of $G\setminus X$ is trivial. As $G$ is 3-connected, every component of $G\setminus X$ has neighbourhood equal to $X$. So $G$ is a $K_{3,m}$ for some $m\geqslant 4$. ∎
1.6. Background on 2-separations

### v3

@@ 4 Nested or crossed: analysing corner diagrams
Suppose that $(C,D)$ is a tri-separation of $G$ that crosses $(A,B)$. If $(C,D)$ is strong, we are done, so we may assume that some vertex $u$ of $G$ of degree three lies in the separator of $(C,D)$. Since $(A,B)$ is a strong tri-separation, the vertex $u$ cannot lie in the centre, so $u$ lies in a link, say it lies in the link for $A$. As a $K_{4}$ has no strong nontrivial tri-separation, $G$ is not a $K_{4}$. By Section 3, $(C,D)$ is nontrivial. Hence we may apply the Nested or crossed: analysing corner diagrams (4) to find that the existence of $u$ implies that all links have size one while the centre consists of a single vertex, and that there are no jumping edges. Since $(C,D)$ is a tri-separation and $u$ has degree three, $u$ has a neighbour $v$ in $C\cap D$. As there are no jumping edges, the neighbour $v$ of $u$ can only lie in the centre. Since $(A,B)$ is a tri-separation, $v$ has a neighbour $w$ in $A$ besides $u$. By symmetry $w\in C$.
Claim 4.0.2.
The vertex $v$ has at least three neighbours in $C$.
Proof of Claim. If the corner $\{B,C\}$ is nonempty, then as $G$ is 3-connected the corner contains a neighbour of $v$, and together with $u$ and $w$ we have found three neighbours of $v$ in $C$. Thus assume that the corner $\{B,C\}$ is empty. By Section 4, at least one adjacent link contains a vertex $y$. Since the corner for $\{B,C\}$ is empty but $y$ has two neighbours either in $B$ or in $C$ (depending on which tri-separator $S(A,B)$ or $S(C,D)$ contains $y$), it follows that $y$ is adjacent to $v$. If $y$ is in the link for $B$, then $u$, $y$ and $w$ are three distinct neighbours of $v$ in $C$, and we were done. So assume that $y$ is in the separator of the strong tri-separation $(A,B)$. Thus $y$ has degree at least four. And since the corner $\{B,C\}$ is empty, the vertex $y$ must have one of its neighbours outside the mixed 3-separator $S(C,D)$ in the corner $\{A,C\}$, and this nonempty corner contains a neighbour of $v$ by 3-connectivity, which is different from $u$ and $y$. $\diamondsuit$
@@
Let $C^{\prime}:=C-u$ and $D^{\prime}:=D$. Then the separator of $(C^{\prime},D^{\prime})$ arises from the separator of $(C,D)$ by replacing the vertex $u$ with the edge $uc$. The only vertex in the separator of $(C,D)$ that might lose a neighbour when moving to $(C^{\prime},D^{\prime})$ is the vertex $v$, which loses its neighbour $u$ in $C^{\prime}$. However, Claim 4.0.2 ensures that $v$ has two neighbours in $C^{\prime}$. Hence $(C^{\prime},D^{\prime})$ is a tri-separation, and it has fewer vertices of degree three in its separator than $(C,D)$. Moreover, $(C^{\prime},D^{\prime})$ crosses $(A,B)$ with the same links and centre as for $(C,D)$, with just one exception: the link for $A$, which consisted of $u$ for $(C,D)$, consists of the edge $uc$ for $(C^{\prime},D^{\prime})$. By iterating at most two times, we obtain a strong tri-separation that crosses $(A,B)$. ∎
@@
(1$\to$(2). We only have to show that $(A,B)$ is strong. This follows from Section 4.
(2$\to$(1). Suppose for a contradiction that $(A,B)$ is crossed by a tri-separation $(C,D)$ of $G$. By Section 4, we may assume that $(C,D)$ is strong. Since $(A,B)$ is strong, $(C,D)$ is nontrivial by Section 3. This contradicts (2). ∎
5 Proof of Applications
Before we prove the (Angry Tri-Separation Theorem)., let us see how it implies Applications. A graph $G$ is essentially 4-connected if it is 3-connected, every nontrivial strong tri-separation has three edges in its separator and the subgraph induced by one side is equal to a triangle, and $G\neq K_{4}$. A graph $G$ is vertex-transitive if the automorphism group of $G$ acts transitively on its vertex set $V(G)$.
@@ 5 Proof of Applications
Let $G$ be a vertex-transitive finite connected graph. We have to show that $G$ either is essentially 4-connected, a cycle, or a complete graph on at most four vertices. By Section 3, $G$ is a cycle, $K_{2}$, $K_{1}$ or 3-connected. Since we are done otherwise, let us assume that $G$ is 3-connected. By the (Angry Tri-Separation Theorem). (1.1), $G$ is internally 4-connected, a $K_{3,m}$ with $m\geqslant 3$, a wheel, or $G$ has a totally-nested nontrivial tri-separation. If $G$ is internally 4-connected, then $G\notin\{K_{4},K_{3,3}\}$ by definition, and all strong tri-separations of $G$ are trivial by Section 3; in particular, $G$ is essentially 4-connected. If $G$ is a $K_{3,m}$ for some $m\geqslant 3$, then $m=3$ since $G$ is vertex-transitive, and $G=K_{3,3}$ is essentially 4-connected since all its strong tri-separations are trivial. If $G$ is a wheel, then $G$ can only be a $K_{4}$ by vertex-transitivity, and $K_{4}$ is a possible outcome.
@@
Proof of Claim. Suppose for a contradiction that there is a vertex $v\in U\cap W$. Since $(U,W)$ is a mixed-separation, there is a vertex $u\in U\setminus W$. Let $\varphi\in\text{Aut}(G)$ send $v$ to $u$. Since $(U,W)$ is a tri-separation and $\varphi$ is an automorphism, the pair $(\varphi(U),\varphi(W))$ is also a tri-separation. Since $(U,W)$ is totally-nested, it is in particular nested with $(\varphi(U),\varphi(W))$. We cannot have $(\varphi(U),\varphi(W))\leqslant(W,U)$ or $(\varphi(W),\varphi(U))\leqslant(W,U)$ since $u\in\varphi(U)\cap\varphi(W)$ and $u\in U\setminus W$. Hence the only two possibilities for $(U,W)$ and $(\varphi(U),\varphi(W))$ to be nested are $(\varphi(U),\varphi(W))\leqslant(U,W)$ or $(\varphi(W),\varphi(U))\leqslant(U,W)$. In either case the inequality is strict since $u$ does not lie in $S(U,W)$ but does so after the application of $\varphi$. This contradicts the choice of $(U,W)$. $\diamondsuit$
Claim 5.0.2.
@@
Proof of Claim. Since $G$ is 3-connected, since $(U,W)$ is nontrivial and since $S(U,W)$ consists of three edges by Claim 5.0.1, it suffices to show that every vertex in $U$ is incident with an edge in $S(U,W)$. The proof is analogous to the proof of Claim 5.0.1. $\diamondsuit$
By Claim 5.0.1 and Claim 5.0.2, every tri-separation in $O$ has three edges in its separator and a side that induces a triangle. As $(A,B)$ was chosen arbitrarily, every totally-nested nontrivial tri-separation has three edges in its separator and a side that induces a triangle. Hence to show that every nontrivial strong tri-separation also has three edges in its separator and a side that induces a triangle, it suffices to show that
Claim 5.0.3.
@@
Proof of Claim. Since $S(U,W)$ consists of three edges (which share no ends by Section 3) and $G[U]=K_{3}$, all vertices in $U$ have degree three. Hence all vertices of $G$ have degree three. Let $(C,D)$ be an arbitrary nontrivial strong tri-separation of $G$. As $(C,D)$ is strong, the separator of $(C,D)$ consists of three edges. Then $(C,D)$ is totally-nested by Section 4. $\diamondsuit$
Combining Claim 5.0.1, Claim 5.0.2 and Claim 5.0.3 yields that $G$ is essentially 4-connected. ∎
@@
6 Understanding nestedness through connectivity
@@
Figure 8: The situation excluded by Section 6
For a depiction of the situation excluded by our next lemma, see Figure 8.
@@
Let $(C,D)$ be a strong tri-separation of $G$ that crosses $(A,B)$ so that $x_{3}$ is in the centre. Since $K_{4}$ has no strong tri-separation, $G$ cannot be a $K_{4}$. By the Nested or crossed: analysing corner diagrams (4) and since $(A,B)$ is half-connected, $x_{3}$ is a vertex and the only element of the centre, all links have size one, and there are no jumping edges. Without loss of generality, $x_{2}$ lies in the link for $C$, and $y\in A\setminus B$. If $y$ lies in the corner for $\{A,C\}$, then we are done. Otherwise, $y$ must lie in the link for $A$. The corner for $\{A,C\}$ must be empty, since otherwise $\{y,x_{3}\}$ would be a 2-separator of $G$, contradicting 3-connectivity. As $y\in S(C,D)$ must have two neighbours in $G[C]$, and since there are no jumping edges, it follows that $yx_{3}$ must be an edge in $G$. But $y\in S(C,D)$ also means that $y$ cannot have degree three as $(C,D)$ is strong, and since $y$ is $x_{3}$-free this means that the edge $yx_{3}$ must not be present in $G$, a contradiction. ∎
Figure 9: The situation in the second and third case of the proof of Section 6
@@
Let us denote the separator of $(A,B)$ by $\{x_{1},x_{2},x_{3}\}$, and let us assume that $x_{3}$ is a vertex and that the separator is externally tri-connected around $x_{3}$. Let us assume for a contradiction that $(A,B)$ is crossed by a strong tri-separation $(C,D)$ so that $x_{3}$ lies in the centre. Since $K_{4}$ has no strong tri-separation, $G$ is not a $K_{4}$. By the Nested or crossed: analysing corner diagrams (4), all links have size one, $x_{3}$ is the only element of the centre, and there are no jumping edges. We distinguish three cases.
@@
($=$) In the third case, $x_{1}$ and $x_{2}$ are edges. By external tri-connectivity, these edges have $x_{3}$-free endvertices $y_{1}$ and $y_{2}$ and there are two internally disjoint paths $P,Q$ from $y_{1}$ to $y_{2}$ avoiding $x_{1},x_{2},x_{3}$. By symmetry assume that the vertex $y_{1}$ lies in the side $A$. Then the two paths $P,Q$ are contained in $G[A]$ and $y_{2}$ is in $A$ as well. Since the link for $A$ has size one, not both of the two paths $P,Q$ can meet it in internal vertices. So $y_{1}$ or $y_{2}$ must lie in the link for $A$. This contradicts Section 6. ∎
@@
Let $(A,B)$ be a tri-separation of $G$ that is externally tri-connected, half-connected, strong and nontrivial. Assume for a contradiction that $(A,B)$ is crossed by a tri-separation $(C,D)$ of $G$. By Section 4, we may assume that $(C,D)$ is strong and nontrivial. As $K_{4}$ has no strong nontrivial tri-separation, $G$ is not a $K_{4}$. Hence by the Nested or crossed: analysing corner diagrams (4), $(A,B)$ and $(C,D)$ cross so that the centre contains a vertex. This contradicts Section 6. ∎
@@
Proof of Section 6.
Note first that $E(A\setminus B,B\setminus A)$ is empty since $G$ is 3-connected. Hence it suffices to find three internally disjoint paths between any pair of vertices in the separator $A\cap B$ avoiding the third vertex, by criterion (:) and Section 6. By assumption, $G\setminus(A\cap B)$ has at least three components, and each component has neighbourhood equal to $A\cap B$ since $G$ is 3-connected. Thus we find three internally disjoint paths for each pair of vertices in $A\cap B$ through these components. ∎
@@
(1). Since every vertex in $X$ has at least one neighbour in every component of $G\setminus X$ and since $B$ includes the vertex-sets of at least two such components, the graph $G[B]$ has minimum degree $\geqslant 2$. From this, we deduce that $B^{\prime}=B$, and so $(A^{\prime},B^{\prime})\leqslant(A,B)$.
@@
(4). If $(A^{\prime},B^{\prime})$ is nontrivial, then $G[A]$ contains a cycle, so two vertices in $X$ are adjacent or $|K|\geqslant 2$. Conversely, suppose now that two vertices in $X$ are adjacent or that $|K|\geqslant 2$. Since $B^{\prime}=B$ and $|B|\geqslant 2$, it follows that $|B^{\prime}|\geqslant 2$. Thus it suffices to show that $A^{\prime}$ contains at least two vertices, by Section 3. If two vertices in $X$ are adjacent, then these two vertices are not reduced to edges in $S(A^{\prime},B^{\prime})$, so they lie in $A^{\prime}$ and we are done. So assume that $|V(K)|\geqslant 2$. Since $V(K)\subseteq A^{\prime}$, we are done as well. ∎
@@
As $(C,D)$ is not half-connected, by 3-connectivity of $G$ its separator $X$ consists of three vertices; and $G\setminus X$ has at least four components. By Section 6, no two vertices in $X$ are adjacent, and every component of $G\setminus X$ is trivial. As $G$ is 3-connected, every component of $G\setminus X$ has neighbourhood equal to $X$. So $G$ is a $K_{3,m}$ for some $m\geqslant 4$. ∎
7 Background on 2-separations